OpenStudy (anonymous):

Find the product. (3p^4)^3 · (p^2)^7

OpenStudy (anonymous):

you multiply exponential of an exponential and add exponentials of base multiplications. 3p^(4x3) . p^(2x7) = 3p^12 . p^14 = 3(p^12+14) = 3 p^26